252 322 6480 0.25 E 4V /12VQuestion 11 1 pts 11. The voltage v across the capacitor fort > 0 2exp(-2t) [2cos(2t) - sin(2t)] V 2exp(-t) [2cos(2t) - sin(2
Answer #1


81 given ckt +=0 M 2 12 12V 0.84 Toast find tol tro fol t20 ( capacitol open inductdk thola) The clt fot to -М w tore (0)2 12

upot) -12 tirot) + tilut) +ilot) = 0 2 4-12 76--260) 2 +i Cot)= -(546) lot) -2 Ampl tallent cot) - Cdw.67) dt Lot) -2 0.25 du

foule turning off collage i te = it cdie dt 2 tit cace too dt 2 ch(di ) +0=0 0.8 do o dt 10 substitute have to find u wl 20 i

0.25 +0.45 +1 -0 (810) 25tus+10=0 doots s=1+2j {=-1-2). Hence Afnz) Et 30 UCHS (A1601 26 + we find Agaq. 410) = 600+)- quolll